OpenStudy (anonymous):

Which is an equation of a circle with center (-5, -7) that passes through the point (0, 0)? A.) (x + 5)2 + (y + 7)2 = 74 B.) (x + 5)2 + (y + 7)2 = 37 C.) (x - 5)2 + (y - 7)2 = 37 D.) (x - 5)2 + (y - 7)2 = 74

OpenStudy (anonymous):

if it passes through origin, that means the distance between (-5, -7) and (0,0) is the radius. the rest is plugging the center of the circle into the equation of a circle

OpenStudy (campbell_st):

well the standard form of a circle is \[(x - h)^2 + (y - k)^2 = r^2\] (h, k) is the centre... so you should be able to match the information... you will need to find the distance from the centre (-5, -7) to the point (0,0) to find the radius.... hope it helps

OpenStudy (anonymous):

8.6 is the distance

OpenStudy (campbell_st):

so what is 8.6^2...?

OpenStudy (anonymous):

73.96

OpenStudy (campbell_st):

so which answer contains your centre and radius

OpenStudy (anonymous):

this are my answer choices : A.) (x + 5)2 + (y + 7)2 = 74 B.) (x + 5)2 + (y + 7)2 = 37 C.) (x - 5)2 + (y - 7)2 = 37 D.) (x - 5)2 + (y - 7)2 = 74

OpenStudy (anonymous):

and im thinking its either A or D

OpenStudy (campbell_st):

ok... so the centre is (-5, -7) and the general form is \[(x - h)^2 + (y - k)^2 = r^2\] (h, k) is the centre and r is the radius... so when you substitute your radius it will be 74 so which equation A or D has the correct centre information..?

OpenStudy (anonymous):

D ?

OpenStudy (campbell_st):

close.... no its A subsitute -5 and its (x - (-5))^2 = (x + 5)^2

OpenStudy (anonymous):

thanks :)